### v1

# X-ray Polarization Detection of the Pulsar Wind Nebula in G21.50.9 with IXPE

Abstract
We present the X-ray polarization observation of G21.50.9, a young Galactic supernova remnant (SNR), conducted with the Imaging X-ray Polarimetry Explorer (IXPE) in October 2023, with a total livetime of approximately 837 ks. Using different analysis methods, such as a space-integrated study of the entire region of the PWN and a space-resolved polarization map, we detect significant polarization from the pulsar wind nebula (PWN) at the center of the SNR, with an average polarization degree of oriented at (north through east). No significant energy-dependent variation in polarization is observed across the IXPE band (2–8 keV). The polarization map, corrected for the effect of polarization leakage, reveals a consistent pattern in both degree and angle, with little change across the nebula. Our findings indicate the presence of a highly polarized central torus, suggesting low levels of turbulence at particle acceleration sites. Unlike Vela, but similar to the Crab Nebula, we observe substantial differences between radio and X-ray polarization maps. This suggests a clear separation in energy of the emitting particle populations and hints at an important, yet poorly understood, role of instabilities in the turbulence dynamics of PWNe.
show]niccolo.dilalla@stanford.edu
1 Introduction
Pulsars are strongly magnetized and rapidly rotating neutron stars, born from the explosive deaths of massive stars, first discovered by Jocelyn Bell Burnell in 1967 using the radio telescope, Interplanetary Scintillation Array in Cambridge, UK (Hewish et al., 1968). As pulsars spin around their axes, they unleash powerful beams of radiation and a persistent powerful stream of relativistic particles into space. These emitted particles, which together with the Poynting flux form the so-called pulsar wind, are accelerated to nearly the speed of light by the intense magnetic and electric fields present in the magnetosphere. As they propagate along the magnetic field lines, they interact with the surrounding interstellar medium or supernova ejecta. This interaction creates shock fronts and gives rise to an extended and complex synchrotron-emitting nebula, from radio to X-ray and MeV energies, called pulsar wind nebula (PWN) (Gaensler & Slane, 2006). PWNe constitute an incredibly fascinating laboratory for exploring fundamental physics in conditions not replicable on Earth, providing valuable insights into relativistic shocks and particle acceleration processes governing the dynamics of relativistic outflows and their influence on the surrounding interstellar environment.
G21.50.9 is a young, composite plerionic supernova remnant (SNR) characterized in X-rays by the presence of a central, bright, 40″ radius PWN surrounded by a much dimmer and softer SNR shell (Slane et al., 2000; Safi-Harb et al., 2001). Deep Chandra and XMM-Newton observations have revealed that the SNR blast wave extends up to 150″ and displays a limb brightening feature at the eastern boundary and knots of enhanced soft X-ray emission above the PWN in the northern direction (Matheson & Safi-Harb, 2005; Bocchino et al., 2005; Matheson & Safi-Harb, 2010; Guest et al., 2019). Observations of G21.50.9 across various electromagnetic wavelengths, from radio to -rays, have unveiled a wealth of information about this source. Using the Very Large Array (VLA) to measure its expansion speed, Bietenholz & Bartel (2008) estimated its age to be yr, making G21.50.9 one of the youngest PWNe known in our galaxy. Its distance to Earth was measured by several authors using different techniques and was found to be around 5 kpc (Safi-Harb et al., 2001; Camilo et al., 2006; Bietenholz & Bartel, 2008; Tian & Leahy, 2008). The first maps at radio wavelengths go back to the 1970–80s (Wilson & Weiler, 1976; Becker & Kundu, 1976; Furst et al., 1988), but even more recent deep searches have failed to detect a radio counterpart of the X-ray SNR shell, with the exception of the northern knot (Bietenholz et al., 2011).
At the heart of G21.50.9 lies the pulsar PSR J18331034, first independently discovered in radio by Gupta et al. (2005) and Camilo et al. (2006), who found a compact pulsating source at the center of the PWN (R.A. = , Dec. = (J2000.0)) with period ms, , surface magnetic field of G, characteristic age of 4.8 kyr, and a spin-down luminosity of . The pulsar remains currently undetected at X-ray energies, while -ray pulsation has been detected with the Fermi Large Area Telescope (Abdo et al., 2013; Smith et al., 2023). Past X-ray observations have only led to upper limits on the pulsating fraction, while still detecting a central compact source surrounded by an elliptical emission region (Slane et al., 2000; Safi-Harb et al., 2001; Camilo et al., 2006). The projected spin-axis direction of PSR J18331034, inferred by fitting the pulsar wind torus using Chandra data, was determined to be at a position angle of north through east (Ng & Romani, 2008).
Polarization observations can play a crucial role in unraveling the complex structure of PWNe like G21.50.9, by offering unique insights into the physical processes at play within these astrophysical sources. Polarimetry, in fact, provides essential information about magnetic fields, particle acceleration mechanisms, and the overall dynamics of these systems, contributing to a more comprehensive understanding of the intricate interplay between pulsars and their cosmic surroundings.
The initial radio polarization maps of G21.50.9 were obtained by Becker & Szymkowiak (1981) with the VLA and by Furst et al. (1988) using the NRO Millimeter-wave Array and revealed a highly ordered, linearly polarized emission distributed in a ring structure, indicating a radial magnetic field configuration. The level of polarization was found to have a minimum towards the center, where little polarization was detected, and to increase with radius up to 20–30% near the boundary of the PWN. Infrared polarimetric observations performed with the Very Large Telescope on the compact core of the PWN revealed a highly polarized emission, with an average value of the linear polarization degree of , and a swing of the polarization angle across the inner nebula consistent with a toroidal magnetic field (Zajczyk et al., 2012). On the other hand, a recent polarimetric study by Lai et al. (2022), using archival high-resolution radio observations taken with VLA, confirmed the global radial magnetic field structure of the PWN and ruled out the existence of any strong large-scale toroidal field components extending beyond the inner nebula.
Unlike Crab, Vela, or MSH 1552 (Bucciantini et al., 2023b; Xie et al., 2022; Romani et al., 2023), which in X-rays all display a morphology dominated by a jet-torus structure, G21.50.9 has a much more amorphous appearance, the torus is barely evident, and there is no sign of a jet. Moreover, among the PWNe that can be resolved by IXPE, G21.50.9 is the only one that shows a tangential polarization pattern in radio. In this regard, this source represents a system quite different from the ones studied previously, and X-ray polarimetry could provide insight to better understand such a difference.
This paper is structured as follows. In Sect. 2 we present the IXPE observation of G21.50.9, together with the initial processing done on the data. Sect. 3 is dedicated to the description of the different types of polarization analyses performed and to the presentation of the main results. The implications and possible physical interpretation of these results are discussed in Sect. 4. Finally, Appendices A and B contain, respectively, additional plots of the polarization analysis and a review of the methods used to evaluate or subtract the effect of polarization leakage.
2 Observations
The Imaging X-ray Polarimetry Explorer (IXPE), successfully launched on December 9, 2021, is the first space observatory entirely dedicated to imaging polarimetry in X-rays (Weisskopf et al., 2022). Result of an international collaboration between NASA and the Italian Space Agency (ASI), IXPE is provided with three identical but independent telescopes, each equipped with a Gas Pixel Detector (GPD) at its focal plane (Bellazzini et al., 2006; Baldini et al., 2021), housed inside a detector unit (DU).
IXPE observed G21.50.9 in three separate segments close in time and combined in a single observation with a total livetime of ks: the first segment lasted from October 5, 2023 at 00:05:39.233 UTC to October 9, 2023 at 02:16:33.292 UTC, the second one from October 10, 2023 at 22:45:54.503 UTC to October 17, 2023 at 16:42:55.034 UTC and the last one from October 19, 2023 at 21:35:44.584 UTC to October 27, 2023 at 15:39:00.534 UTC. IXPE Level-2 data, publicly available for download on the HEASARC archive111https://heasarc.gsfc.nasa.gov/docs/ixpe/archive/ (ObsID 02001199), were reprocessed and analyzed according to the standard technique using the version 20240125 of the IXPE Instrument Response Functions (IRFs) released on the CALDB222https://heasarc.gsfc.nasa.gov/docs/ixpe/caldb/ on February 28, 2024 (validity date 20230702). Before analyzing the data, we performed an initial background rejection following the procedure described in Di Marco et al. (2023) to reduce the instrumental background due to cosmic-ray particles interacting in the GPD by approximately 40%. We also removed the short time intervals where sudden spikes in the counting rate are observed connected to solar activity and the South Atlantic Anomaly boundary, and corrected the absolute sky coordinates by adjusting the WCS keywords in the FITS header to align the reference frame to Chandra (TCRPX7 = 304, TCRPX8 = 301.5). These are the data files used from now on in this paper. Fig. 1 shows the count map resulting from the combination of the three IXPE telescopes over the full energy range (2–8 keV). The white circular region (radius 0.8′ centered on the pulsar location) is used to select the inner PWN, while the yellow annulus (inner/outer radii of 3′ and 4′ from the pulsar) defines the background extraction region used for the space-integrated analysis described in Sect. 3.1.
G21.50.9 is observed with Chandra routinely as a calibration source. Overlaid in green in Fig. 1 are the contours of a deep image created by merging 18 individual observations listed in Table 1 (total exposure ks), all of which had the remnant placed in the ACIS-S detector, at the default aimpoint of the S3 CCD chip. The events were combined with the merge_obs script in the ciao333https://cxc.cfa.harvard.edu/ciao/ analysis software. A count map was then created using events in the 0.5–8.0 keV energy band, and adaptively smoothed with the dmimgadapt tool using a cone filter with minimum/maximum smoothing radii of 0.5″ and 150″ spanning 300 logarithmically spaced scales, with a minimum of 75 counts within the smoothing kernel. The resulting map is used as background image in all polarization maps shown in the paper (e.g. Fig. 3).
ObsID | Exposure (ks) | ObsID | Exposure (ks) |
159 | 14.85 | 3700 | 9.54 |
1230 | 14.56 | 5166 | 10.02 |
1433 | 14.97 | 5159 | 9.83 |
1717 | 7.54 | 6071 | 9.64 |
1770 | 7.22 | 6741 | 9.83 |
1838 | 7.85 | 8372 | 10.01 |
1554 | 9.06 | 10646 | 9.54 |
2873 | 9.83 | 14263 | 9.57 |
3693 | 9.78 | 16420 | 9.57 |
3 Polarization Analysis
In this section, we describe the different analysis methods used to measure the polarization of G21.50.9. We performed two main types of analysis: an integrated study of the entire region of the PWN (Sect. 3.1) and a space-resolved measurement aimed at mapping the polarization pattern across the nebula (Sect. 3.2).
3.1 Space-integrated analysis
As is typically done for other IXPE observations (Bucciantini et al., 2023b; Xie et al., 2022; Negro et al., 2023), using the spatial regions described at the end of Sect. 2 and shown in Fig. 1, we studied the polarization of the entire PWN in two different ways: a model-independent polarimetric analysis using ixpeobssim (Baldini et al., 2022) and a spectro-polarimetric fit with the Multi-Mission Maximum Likelihood (3ML) framework (Vianello et al., 2015). For the former, we utilized the xpselect and xpbin applications of ixpeobssim (v31.0.1) to first select photons belonging to the source and background regions and then bin these filtered data files using the PCUBE algorithm (unweighted analysis). This binning routine combines the event-by-event Stokes parameters, subtracts the background contribution, and computes the degree (PD) and angle (PA) of polarization of the source, along with the associated errors, according to the procedure of Kislat et al. (2015). PA represents the electric vector position angle and is conventionally defined anticlockwise relative to the local north (towards northeast). Combining the data from the three IXPE DUs over the 2–8 keV energy range, we obtained with this method a PD for the PWN of oriented at . This PA is similar, although in tension at 3 level, with the measured spin-axis direction of PSR J18331034. We also repeated the same analysis by dividing the data into two energy bins, 2–4 keV and 4–8 keV, and obtained a result that is consistent within uncertainties with a constant polarization throughout the IXPE energy range. The results are summarized on the left side of Table 2 and visualized in Fig. 2.
Parameter | Energy range | Value | Parameter | Value |
2–8 keV | (fixed) | |||
PD | 2–4 keV | |||
4–8 keV | PD | |||
2–8 keV | PA | |||
PA | 2–4 keV | Q/I | ||
4–8 keV | U/I |
Note. — Summary table of the PCUBE analysis (left) and spectro-polarimetric fit (right) of the space-integrated PWN. Errors are statistical only at 1 level. The spectro-polarimetric fit is performed using 3ML in the 2–8 keV energy range assuming Gaussian statistics and resulted in a reduced chi-square value of (-value = 0.02).
For the spectro-polarimetric analysis, instead, we started from the same filtered data, produced earlier for the source and background regions, and used the xpbin application (algorithms PHA1, PHA1Q and PHA1U, unweighted analysis) to compute the Stokes I, Q, and U spectra for the three DUs. These files can then be jointly fit in 3ML or xspec444https://heasarc.gsfc.nasa.gov/xanadu/xspec/ to simultaneously recover the spectral and polarimetric parameters of the user-defined models. For G21.50.9, we modeled the observed spectrum with an absorbed power-law, with absorption given by the TbAbs555https://astromodels.readthedocs.io/en/latest/notebooks/TbAbs.html model with wilm abundances (Wilms et al., 2000). According to what was found by Guest et al. (2019), the column density parameter was fixed to the value . Regarding the polarimetric model, based on the result of the previous PCUBE analysis, we chose to employ a model with a constant degree and angle of polarization. To account for potential cross-calibration uncertainties among the different IXPE telescopes, we also added and left free to vary a multiplicative normalization constant for DU2 and DU3 relative to DU1. The best-fit results obtained using 3ML are provided on the right side of Table 2. The power-law index we find is , which is somewhat larger than expectations based on the work by Guest et al. (2019), most likely due to contamination from the much softer external halo and to a different size of the source region. Please note that the uncertainty quoted here represents the statistical error only and does not include any systematics due to potential calibration issues (see Appendix A for further discussion). The polarimetric results agree well with the PCUBE analysis with a best-fit PD = and PA = . Fig. 2 gives a summary of the main results of the space-integrated polarimetric analysis of the PWN.
3.2 Space-resolved analysis
To map the polarization of G21.50.9 in the PWN region, we again used the xpbin tool of ixpeobssim, this time with the PMAP algorithm. This routine processes the provided data files by binning the Stokes parameters I, Q and U in sky coordinates, thus allowing us to measure the polarization pattern across the field of view. In particular, we binned the data using a grid of squared pixels, each with a side of 0.1944′, considering photons with energies in the full IXPE energy band (2–8 keV). Then, we used the implemented methods to perform a two-dimensional convolution of the original Stokes maps with a unit-filled kernel matrix, assigning to each pixel the sum of the contents of itself and its eight neighbors. In this way, adjacent pixels become highly correlated and the effective size of the pixel triples and becomes comparable to the IXPE angular resolution (half-power diameter 25–30″), while still maintaining the original binning grid.
It has been well known since the beginning of the IXPE mission that any imperfections in the reconstruction of the photon absorption points inside the GPD, due to the event reconstruction algorithm (Baldini et al., 2021) or the spatial resolution of the GPD itself, can cause the appearance of a radially polarized halo around the source, commonly referred to as polarization leakage (for a detailed overview of the effect, see Bucciantini et al. (2023a)). Being an intrinsic effect of the reconstruction algorithm or the GPD, this phenomenon is present in all IXPE observations, but in practice it is only really relevant for the analysis of a handful of bright extended sources, which exhibit a very sharp intensity gradient. Since the polarization pattern induced by the leakage is radial, its contribution effectively averages out when selecting a circular region that includes the entire source, as done in Sect. 3.1. For the space-resolved polarization analysis, instead, polarization leakage can strongly affect the results, and a dedicated analysis is required to estimate and subtract its contribution.
Given the morphology of G21.50.9, and its small relative size comparable to the IXPE angular resolution, it is expected that the effect of polarization leakage will be particularly relevant in the outskirts of the PWN. Its contribution was evaluated using different techniques, including a full ixpesim/ixpeobssim (Di Lalla, 2019) simulation of the source using GEANT4 (Agostinelli et al., 2003), the Mueller matrix formalism described in Bucciantini et al. (2023a), a hybrid machine learning and analytic track reconstruction method (Cibrario et al., 2023) and, finally, the recently released LeakageLib package (Dinsmore & Romani, 2024). The latter tool, in particular, implements the most accurate model of the polarization leakage to date using in-flight calibrated 2D PSFs, and its routines allow the user to predict the leakage pattern based on a Chandra image and subtract it from the measured IXPE polarization map. A detailed summary of the different approaches can be found in Appendix B.
Figure 3 shows the detected IXPE polarization pattern, overlaid on top of the deep Chandra count map of G21.50.9 zoomed over the PWN region. The effect of polarization leakage has been estimated and subtracted with LeakageLib, using the same Chandra image to model the source. The measured polarization information is encoded in a set of segments, with a length proportional to the measured PD, oriented according to the PA (measured east of north) and with a color that depends on the level of significance of the detection: green lines represent measurements with significance , while the black color corresponds to significance between and . The white lines represent the polarization pattern measured in radio (Lai et al., 2022). Measurements consistent with no polarization within are not shown. As can be seen in Fig. 3, the detected pattern in the central region of the PWN, where the detection is most statistically significant, is roughly consistent with a constant PD of 11–12% at an angle of 30–35°. The detection in the outer regions has generally lower significance and the measured PD and PA values vary slightly across the nebula. This fact explains why the space-integrated analysis (Sect. 3.1), which encompasses the whole PWN, measured a lower degree of polarization. The only notable exception is the NW region (top right) of the PWN, where the degree of polarization becomes progressively higher up to a PD of 20–30% and an orientation around 20–25°. No significant polarization was detected outside the PWN.
4 Discussion
Polarization studies provide a powerful diagnostic tool for probing the magnetic field structure and particle acceleration mechanisms in PWNe. In the case of G21.50.9, IXPE observations allow us to explore the polarization properties of the X-ray-emitting electrons and compare them with those observed in the radio band. This comparison offers crucial insights into the evolution of the PWN’s magnetic field and its interaction with the surrounding SNR environment.
The model assuming the same outer polarization structure as seen in radio—where the magnetic field is predominantly radial with respect to the pulsar—predicts that in an annulus with inner and outer radii of ″ and ″, respectively, the PA should be tangentially directed, implying an inferred radial magnetic field, with a background-subtracted and leakage-corrected PD . However, our IXPE analysis finds no significant polarization in this region, strongly suggesting that the magnetic field structures traced by the radio- and X-ray-emitting electrons are different. In this respect, G21.50.9 appears more similar to the Crab Nebula than to Vela, where the radio and X-ray polarization structures are well aligned (Xie et al., 2022; Liu et al., 2023). Interestingly, both G21.50.9 and the Crab Nebula are believed to be in the free-expansion phase, where the PWN expands into the cold SNR ejecta (Bucciantini et al., 2003), and strong Rayleigh-Taylor instabilities shape the outer regions (Jun, 1998; Bucciantini et al., 2004). Vela, on the other hand, is likely in a later evolutionary phase, where interaction with the SNR reverse shock is modifying the structure and dynamics of the nebula (van der Swaluw et al., 2003).
Given that G21.50.9 is at the resolution limit of IXPE, we cannot directly determine which specific nebular structures contribute most to the observed polarization. In particular, two competing scenarios remain indistinguishable: (i) a highly polarized central torus dominating the polarized emission of the source, as suggested by infrared observations (Zajczyk et al., 2012), or (ii) a bulk PWN magnetic field that is more uniform but with a lower degree of polarization (). However, the latter scenario is inconsistent with radio observations, which show a clearly radial magnetic field pattern (Lai et al., 2022).
Nebula Model | Torus PD | Torus PA |
Unpolarized | ||
Radio-like polarized |
Note. — Different model configurations for the inner torus and nebula that reproduce the polarization properties measured by IXPE in the central region of the PWN. The intrinsic polarization values for the torus are obtained through model fitting using ixpeobssim simulations for each nebula configuration.
To evaluate the possible contribution of a highly polarized torus, we simulated two different scenarios using ixpeobssim: (i) a uniformly polarized torus embedded in an unpolarized nebula, and (ii) the same torus within a nebula that follows the polarization structure and degree observed in the radio band at 5 GHz (Lai et al., 2022). In both cases, the polarized torus is modeled as an elliptical region (″″) with uniform polarization, inclined by and centered on the Chandra X-ray intensity peak. We varied the torus PD and PA in the simulations until we achieved consistency with the polarization properties measured in the central region of the PWN (space-integrated PCUBE analysis within a radius ″ from the PSR: PD , PA ). This region was chosen because it provides the highest polarization significance while minimizing leakage effects, which are not simulated by ixpeobssim (see Appendix B.2). The results of these simulations for the two model configurations are shown in Table 3. Both models yield large uncertainties in the intrinsic polarization properties of the torus, and neither scenario is strongly favored, though the unpolarized nebula case appears somewhat closer to the infrared observations. In either case, the results are consistent with previous IXPE studies of other PWNe (Bucciantini et al., 2023b; Romani et al., 2023; Xie et al., 2022), suggesting that even in the case of G21.50.9, close to the termination shock—where particles are likely being accelerated to high energies—turbulence remains moderate ( at most ), indicating a relatively ordered magnetic field.
Our findings suggest that a two-zone model for G21.50.9 is probably preferred: a central bright X-ray torus dominated by a globally toroidal magnetic field, surrounded by an outer layer affected by Rayleigh-Taylor instabilities, similar to the filamentary network seen in the Crab Nebula, which stretches the magnetic field radially. However, the overall low integrated polarization of G21.50.9 distinguishes this source from other well-studied PWNe like the Crab, Vela, and MSH 1552. This fact, along with differences in the radio polarization pattern and the absence of a strong jet-torus structure, suggests that X-ray-emitting particles may populate a larger region of the PWN where turbulence is stronger or where magnetic field geometry varies significantly due to internal dynamics or enhanced diffusion.
Appendix A Spectro-polarimetric analysis
In this section of the Appendix, we provide additional plots to support the spectro-polarimetric analysis of the entire PWN, as described in Sect. 3.1. Fig. 4 presents the Stokes I, Q and U spectra for the three DUs, along with their best-fit models and the residuals from the 3ML analysis. While the Stokes Q and U spectra are well described by the model, the Stokes I spectral fit shows significant deviations at low energies, most likely due to instrumental calibration issues. These discrepancies become more noticeable as the statistical uncertainties approach the level of systematic errors. To assess the potential impact of these issues on the polarization measurement, we repeated the same analysis in two additional configurations: (i) applying the off-axis vignetting correction to the on-axis IXPE effective area (version 20240125, used throughout the paper) using the HEASARC tool ixpecalcarf666https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/help/ixpecalcarf.py.html, and (ii) employing the previous time-independent version of the IXPE IRFs (version 20230526, released on June 16, 2023). In both cases, while the best-fit power-law index varied within the range 1.85–1.98, the resulting PD and PA values remained virtually unchanged. Overall, despite the minor calibration uncertainties that affect the spectral analysis, the stability of the polarization measurements confirms the robustness and reliability of our polarimetric results.
Appendix B Polarization leakage evaluation and correction
Fig. 5 shows the uncorrected polarization map of G21.50.9 resulting from the space-resolved analysis (2–8 keV). Compared to Fig. 3, an outflow pattern is clearly visible, caused by the superposition of the intrinsic, nearly uniform polarization of the source and the radial footprint due to leakage. The purpose of this appendix is to review and compare the various methods used to evaluate or subtract the effect of the polarization leakage. As reported in Sect. 3.2, we employed four different techniques: the LeakageLib software package, a full ixpesim/ixpeobssim simulation of the source, a joint machine learning and analytic event reconstruction method, and the Mueller matrix approach, each one described in the following sections. Although each method has some limitations, LeakageLib is currently the most advanced correction technique. For this reason, we used it to generate Fig. 3 and adopted it as the reference for comparison with the other methods (Fig. 6). Overall, all four methods provide consistent results, reinforcing confidence in our polarization leakage correction techniques. The LeakageLib and Mueller matrix methods yield nearly identical outcomes, while the hybrid event reconstruction provides an independent confirmation by directly mitigating leakage effects. The ixpesim/ixpeobssim simulation slightly overestimates leakage correction but remains within acceptable limits. These comparisons highlight the robustness of our polarization analysis and the importance of accurately modeling leakage effects in extended sources like G21.50.9.
B.1 LeakageLib
The LeakageLib (Dinsmore & Romani, 2024) package introduces a new model and algorithm for correcting polarization leakage in IXPE data. Compared to the linearized approach of the Mueller matrix formalism, described later in Sect. B.4, this method allows for a more general correlation between the spatial offset of the reconstructed impact point and the inferred polarization direction. Thanks to the implementation of a 2D displacement model, which accounts for both longitudinal and transverse components of the offset, the formalism implemented in LeakageLib can be applied to a more general model of PSF, not necessarily azimuthally symmetric. To better represent the data, this software uses a set of sky-calibrated 2D PSFs, one for each IXPE telescope, derived from a set of bright, weakly polarized point sources observed by IXPE.
For extended sources, the tool provides routines to estimate the expected Stokes maps (I, Q and U) due to polarization leakage based on a Chandra observation of the source. The Chandra count map is assumed to represent the true morphology given its far superior angular resolution compared to IXPE. However, in the current release (v1.1.0), no PSF dependence on off-axis angle or event energy has yet been implemented. For G21.50.9, we modeled the source using the same Chandra map employed throughout the paper. The output Stokes maps, containing the predicted signal induced by leakage, are then normalized to match the observed data and properly rebinned using the same grid as in Fig. 5. For all IXPE telescopes, the content of the Stokes Q and U maps is finally subtracted from the polarization map obtained in Sect. 3.2.
B.2 ixpesim/ixpeobssim simulation
The ixpeobssim (Baldini et al., 2022) software allows the user to simulate an IXPE observation starting from a generic model of the source or a Chandra observation. However, ixpeobssim alone cannot predict polarization leakage, as this effect is not included in the standard IRFs. To enable leakage prediction, ixpeobssim offers an interface to ixpesim (Di Lalla, 2019), which implements a detailed photon-by-photon GEANT4 simulation of the physical interactions in the GPD. As described in the official documentation777https://ixpeobssim.readthedocs.io/en/latest/ixpesim.html, a list of photons can be simulated in ixpeobssim up to the Beryllium window on top of the GPD using the xpphotonlist application. These photons are then propagated throughout the detector using ixpesim and turned into a list of actual photoelectron tracks for the events that trigger the detector. The output file containing the track images (comparable to the IXPE Level-1 data) can be reconstructed using the same event reconstruction algorithm used for the flight data, and a final step using xpsimfmt, part of ixpeobssim, closes the loop by formatting it in a way virtually identical to the actual IXPE Level-2 data.
Based on a Chandra observation of G21.50.9 (Obs ID: 5166), we ran 100 full ixpesim/ixpeobssim simulations, equivalent to a total exposure time of 40 Ms. We modeled the source as intrinsically unpolarized, so that any pattern detected in the output polarization map could be attributed to the leakage effect. By rescaling and rebinning the resulting Stokes maps with the same binning scheme as in Fig. 5, their contribution can be subtracted pixelwise to correct the polarization leakage effect. Fig. 6 shows a detailed comparison between the outcome of LeakageLib and the other three methods. In particular, top row is relative to the ixpesim/ixpeobssim simulation, and from left to right, the three columns show the corrected polarization map obtained with this method (2–8 keV) and the differences in sigmas of the maps of degree and angle of polarization compared to LeakageLib. In contrast to all other methods, this approach appears to over-correct the leakage effect, inducing a spurious tangential pattern. This could be due to an imperfect tuning of the detector parameters in ixpesim, given the observed time evolution of the GPD. Nevertheless, the differences measured between the two methods are within 3, although clearly spatially correlated due to the over-prediction of the leakage amplitude by the simulation.
B.3 Hybrid event reconstruction
The hybrid event reconstruction is a recently developed method, described in detail in Cibrario et al. (2023), which combines a new machine learning-based prediction of the photon impact point with a standard moment analysis reconstruction of the photoelectron track, commonly used for IXPE data. This joint approach exploits the better performance achieved by convolutional neural networks in predicting the conversion points from the track images, improving the IXPE polarization capabilities (expressed by the modulation factor) and partially mitigating the effect of polarization leakage (Cibrario et al. in prep).
We reprocessed the IXPE Level-1 data of this observation, publicly available on the HEASARC archive along with the Level-2 data, using the hybrid event reconstruction method. This technique is particularly valuable for comparison, as it is the only approach that directly mitigates polarization leakage rather than subtracting it. Thus, it provides an independent and complementary assessment. As described in Sect. 3.2, we binned the reprocessed data following the same recipe using the PMAP algorithm of xpbin to produce a new polarization map. The central row of Fig. 6 shows the map resulting from this method and its comparison with LeakageLib. Despite the different methodology, the results in the high-significance region of the PWN are fully consistent with LeakageLib within 1.
B.4 Mueller matrix formalism
The Mueller matrix approach, thoroughly discussed in Bucciantini et al. (2023a), is essentially a generalization of the PSF for polarized observations, where the elements of the Mueller matrix describe how the intrinsic Stokes parameters mix with each other due to the effect of polarization leakage. The primary assumption of this method is that the displacement of the reconstructed absorption point from the true position is correlated with the direction of the reconstructed polarization plane (preferentially displaced along the direction of the polarization plane). This simplifies the problem to a linear displacement model. The matrix elements can be derived either theoretically from a given PSF, or by fitting the expected functional form using in-flight IXPE data. In either case, computing the Stokes maps of any source is reduced to a simple convolution with a Chandra count map. For G21.50.9, we adopted parameters derived from the IXPE observation of Cygnus X-1 (Obs ID: 01002901), one of the brightest, weakly polarized point sources observed by IXPE. Unlike LeakageLib, the tool implementing this formalism assumes a circular, azimuthally symmetric PSF, as the one provided by ixpeobssim. Although this is an approximation of the real PSF (see Dinsmore & Romani (2024) for a detailed comparison), this model still provides a reasonable estimate of the average effect across the three IXPE telescopes. As previously done for LeakageLib, the leakage estimation is based on a Chandra observation of G21.50.9, and the resulting Stokes maps are properly normalized, rebinned and finally subtracted from the uncorrected polarization map, to allow a meaningful comparison. As for the other methods, the subtracted polarization map and its comparison with LeakageLib is reported in Fig. 6 (bottom row). The results of these two methods are in excellent agreement, with differences in the PD and PA maps well below 1, most likely due to differences in PSF parameterization.